Two moles of an ideal gas undergo a reversible isothermal expansion from 2.41×10−2 m3 to 4.42×10−2 m3 at a temperature of 26.7 ∘C.

What is the change in entropy ΔS of the gas?

Express your answer numerically in joules per kelvin.
Answer #1

here,

number of moles , n = 2

temprature , T = 26.7 degree C

T = 299.7 K

Vfinal = 4.42 * 10−2 m3

Vinitial = 2.41 * 10−2 m3

heat exchanged , Q = work done

Q = n * R * T * ln(Vfinal/Vinitial)

Q = 2 * 8.314 * 299.7 * ln(4.42/2.41)

Q = 3022.5 J

change in entropy = Q / T

change in entropy = 3022.5 /299.7

change in entropy of the gas is 10.09 J/K